OpenStudy (anonymous):

find linearization of (64+3x^2)^(-1/2)

OpenStudy (anonymous):

this means find the equation of the line tangent to the curve but to do this you need a point as well

OpenStudy (anonymous):

at a=0 that would be the point right?

OpenStudy (anonymous):

i plug zero into the derivative to get the slope?

OpenStudy (anonymous):

i don't know, you cannot assume the point is a = 0 you must be told

OpenStudy (anonymous):

you can only find the linearization at a point, so you have to be told both the function and a point

OpenStudy (anonymous):

that is the point i was given i just forgot to write it

OpenStudy (anonymous):

would this be correct.. y=1/8+512(x-0)?

OpenStudy (anonymous):

oh ok find the derivative, replace x by 0 and get the slope, then use point slope formula lets see what we get

OpenStudy (anonymous):

\[f(x)=(64+3x^2)^{-1/2}\] \[f'(x)=-\frac{3x}{(64+3x^2)^{\frac{3}{2}}}\]

OpenStudy (anonymous):

hmm i get \[f'(0)=0\] did i make a mistake?

OpenStudy (anonymous):

oh no i got zero too

OpenStudy (anonymous):

so plug that into point slope formula is that what i do?